It takes 3 hours to print a local news paper using four printers.
(1)How long will it take six printers to print the local news paper?
(2)How many printers are needed if the printing of the paper must be completed in four hours?

Answer by richwmiller(17219)

3*4=12 printer hours
1) 12/6=2 hours
2) 12/4= 3 printers